OpenStudy (anonymous):

solve for the negative value of x: |2x+5|+1=13

jimthompson5910 (jim_thompson5910):

|2x+5|+1=13 |2x+5|=13-1 |2x+5|=12 2x+5=12 or 2x+5=-12 2x=12-5 or 2x=-12-5 2x=7 or 2x=-17 x = 7/2 or x = -17/2 So the answer is x = -17/2